OpenStudy (anonymous):

sketch the graph y=-3tanpiex

OpenStudy (anonymous):

sketch the graph of the function( include two full period and 2 consecutive asymptotes of the graph ) use a graphing utility to verify your results. A)y= -3tanpiex B)y=2sec4x C)y=cscx/3 d)y=3cotpiex/2

OpenStudy (anonymous):

To sketch a graph of\[y=\tan (\pi x)\] You first need to find the period. The period of a tangent function is \[\pi/B\] Your B value in this equation is \[\pi\]

OpenStudy (anonymous):

So your period is \[\pi/\pi=1\]Next you divide the period in to four equal sections. So 1 divided by 4 = 1/4 So your x values will be: x=-4/4; x=-3/4; x=-2/4; x=-1/4; x=0;x= 1/4; x=2/4; x=3/4; x=4/4 Now for each x value substitute in to the equation.\[y=-3\tan (\pi/4)\] For example, x=0, your y value will be 0. So your point will be (0,0) For x=-1/4, the y value will be 3. So your point will be (-1/4,3)

OpenStudy (anonymous):

For the equation\[y=2\sec 4x\]Again you find the period of the function.l For the sec function the period is \[2\pi/B\] Here your B value is 4. So the period is \[\pi/2\]Again divide the period into four equal sections. So your x values will be \[x=-4\pi/8; x=-3\pi/8; x=-2\pi/8; x=-\pi/8; x=0; x=\pi/8; x=2\pi/8; x=3\pi/8; x=4\pi/8\] Again substitute in to the equation and solve for y.